# Contrasting GHZ and W-state Entanglement Dynamics due to Correlated Markov Noise

### v1

@@ 1 Introduction
Entanglement is a fundamental feature of quantum mechanics with no classical analogue [1]. It serves as a central resource in quantum information processing, enabling protocols in quantum computing [2], quantum communication [3], and quantum cryptography [4]. However, in realistic scenarios, quantum systems are inevitably coupled to their environment, resulting in decoherence [5] and the degradation of entanglement [6, 7, 8, 9]. Understanding the dynamics of entanglement in open quantum systems is therefore essential for the development of scalable quantum technologies [10, 11].
Previous theoretical studies have demonstrated the potential for entanglement modulation and rebirth [12, 13, 14] as well as the effects of noise correlations on qubit dynamics [15, 16]. In this work, we examine the dynamics of multipartite entanglement in a system of three uncoupled qubits, each subject to local, classical environmental noise. By introducing statistical cross-correlations between the noise sources, we explore how such correlations can mitigate or accelerate entanglement decay. Using tripartite negativity as our entanglement measure, we analyze the time evolution of two fundamental three-qubit states, the GHZ state and the W state. Our results demonstrate distinct responses by these states to correlated versus anticorrelated noise, revealing conditions under which multipartite entanglement can be sustained despite environmental interactions.
@@ 5 Conclusion
We have investigated the dynamics of three-qubit entanglement under semiclassical correlated noise, analyzing both dephasing and amplitude noise using a Markovian master equation approach. By introducing tunable statistical correlations between local noise environments, we identified distinct mechanisms through which such correlations influence multipartite entanglement.
Under dephasing noise, GHZ and W-type entanglement respond oppositely to the sign of noise correlations. The GHZ state, whose entanglement is intrinsically global, benefits from anticorrelated environments that can, in the ideal limit, completely suppress decoherence at the point $S(-\tfrac{1}{2},-\tfrac{1}{2},-\tfrac{1}{2})$. In contrast, the W state, which retains bipartite correlations in its reduced subsystems, is stabilized by correlated environments, achieving full preservation at $Q(1,1,1)$. These contrasting dependencies arise from the additive versus pairwise (product-like) roles of the correlation parameters in the respective tripartite negativities, giving us a direct view into their fundamentally different underlying entanglement structures.
@@
Together, these results elucidate how semiclassical noise correlations influence and, in certain regimes, can stabilize multipartite entanglement. Beyond their theoretical relevance, this framework provides intuition for how different entanglement structures respond to correlated environments and may guide future considerations for noise-prone multi-qubit quantum architectures.

### v2

@@ 1 Introduction
Entanglement is a fundamental feature of quantum mechanics with no classical analogue [1]. It serves as a central resource in quantum information processing, enabling protocols in quantum computing [2], quantum communication [3], and quantum cryptography [4]. However, in realized scenarios, quantum systems are inevitably coupled to their environment, resulting in decoherence [5, 6, 7] and the degradation of entanglement [8, 9, 10, 11, 12]. Understanding the dynamics of entanglement in open quantum systems is therefore essential for the development of scalable quantum technologies [13, 14].
Previous theoretical studies have demonstrated the potential for entanglement modulation and rebirth [15, 16, 17] as well as the effects of noise correlations on qubit dynamics [18, 19]. In this work, we examine the dynamics of multipartite entanglement in a system of three uncoupled qubits, each subject to local, classical environmental noise, known to inevitably lead to disentanglement [20, 21, 22]. By introducing statistical cross-correlations between the noise sources, we explore how such correlations can mitigate or accelerate entanglement decay. Using tripartite negativity as our entanglement measure, we analyze the time evolution of two fundamental three-qubit states, the GHZ state and the W state. Our results demonstrate distinct responses by these states to correlated versus anticorrelated noise, revealing conditions under which multipartite entanglement can be sustained despite environmental interactions.
@@ 5 Conclusion
In summary, we have investigated the dynamics of three-qubit entanglement under correlated noise, analyzing both dephasing and amplitude noise using a Markovian master equation approach. By introducing statistical correlations between pairs of local noise environments, we identified distinct mechanisms through which such correlations influence multipartite entanglement for different initial states.
Under dephasing noise, GHZ and W-type entanglement respond oppositely to the direction of noise correlations. The GHZ state, whose entanglement is intrinsically global, benefits from anticorrelated environments that can, in the ideal limit, completely suppress decoherence at the point $S(-\tfrac{1}{2},-\tfrac{1}{2},-\tfrac{1}{2})$. In contrast, the W state, which retains bipartite correlations in its reduced subsystems, is stabilized by correlated environments, achieving full preservation at $Q(1,1,1)$. These contrasting dependencies arise from the additive versus pairwise (product-like) roles of the correlation parameters in the respective tripartite negativities, giving us a direct view into their fundamentally different underlying entanglement structures.
@@
Together, these results elucidate how noise correlations can influence and, in certain regimes, stabilize multipartite entanglement. Beyond their theoretical relevance, this framework provides intuition for how different entanglement structures respond to correlated environments and may guide future considerations for noise-prone multi-qubit quantum architectures.